OpenStudy (anonymous):

The circle is tangent to the line 4x+3y = 4 at point (4 , -4) and the center is on x-y = 7. Find the equation of the circle.

OpenStudy (anonymous):

calculate the perpendicular distance between line tangent and the center it should work

OpenStudy (akshay_budhkar):

even i can say that.. plz solve it superhero

OpenStudy (anonymous):

lol but i don't wanna do it please dumbcow is here he is good better then me he will figure it out

OpenStudy (akshay_budhkar):

guys we will scare kris away from open study!!!! plz help him i am not wanting to solve it coz i am making silly mistakes.. (lol)

OpenStudy (dumbcow):

haha ok first determine the equation of tangent line in slope-intercept form \[y=-\frac{4}{3}x +\frac{4}{3}\] Now the line perpendicular to this line must go through the center of circle. slope is opposite reciprocal of -4/3 \[m = \frac{3}{4}\] goes through point (4,-4) \[-4 = \frac{3}{4}(4) + b\] \[b = -7\] Perpendicular line:\[y = \frac{3}{4}x - 7\] Now we also know the center is on the line\[y = x-7\] Therefore the center is the point where they intersect: x-7 = (3/4)x - 7 x = 0, y =-7 Finally, the radius is the distance from (0,-7) to (4,-4) \[r^{2} = (4-0)^{2} +(-4-(-7))^{2} = 4^{2}+3^{2} = 25\] Equation of the circle: \[x^{2}+(y+7)^{2} =25\]

OpenStudy (anonymous):

Excellent Solution :D